# A Library for Learning Neural Operators

### v4

@@ Abstract
We present NeuralOperator, an open-source Python library for operator learning. Neural operators generalize neural networks to maps between function spaces instead of finite-dimensional Euclidean spaces. They can be trained and inferenced on input and output functions given at various discretizations, satisfying a discretization convergence properties. Built on top of PyTorch, NeuralOperator provides all the tools for training and deploying neural operator models, as well as developing new ones, in a high-quality, tested, open-source package. It combines cutting-edge models and customizability with a gentle learning curve and simple user interface for newcomers.
@@ 1 Introduction
Deep neural networks have been considered to accelerate the solution of PDEs by mapping from parameters directly to the solution on a given discretization (Guo et al., 2016; Bhatnagar et al., 2019; Wang et al., 2020; Gupta and Brandstetter, 2023). However, they can only learn mappings between finite-dimensional spaces, not function spaces. In other words, the solutions learned are tied to a fixed discretization. In particular, there is no guarantee that neural networks generalize to other discretizations, and they often perform poorly when interpolated to higher resolutions (Azizzadenesheli et al., 2024).
To address these limitations, a new class of machine learning models, known as neural operators, was proposed (Li et al., 2020, 2021; Kovachki et al., 2023). While standard neural networks learn mappings between fixed-size discretizations, i.e., finite-dimensional spaces, neural operators can directly learn mappings between functions, i.e., infinite-dimensional spaces (Kovachki et al., 2024; Azizzadenesheli et al., 2024). They are built from first principles to ensure a discretization convergence property: a neural operator, with a fixed set of parameters, can be applied to input functions given at any discretization. Specifically, for an input function given at various discretizations, the outputs differ only by a discretization error that converges to zero as the discretization is refined.
@@ 2 The NeuralOperator Library
NeuralOperator is open-sourced under MIT license [^1] Neural Operator is hosted at https://github.com/neuraloperator/neuraloperator.. In the following, we provide details on its design principles and functionalities; see Figure 1 for an overview.
The NeuralOperator library builds on the following guiding principles:
  • Resolution-agnostic: As the crucial difference to existing frameworks, modules in NeuralOperator, such as data loaders, architectures, and loss functions, should be applicable to functions at various discretization. As an example, our implementation of the Fourier Neural Operator (FNO) (Li et al., 2021), just like the theoretical model, can take inputs on regular grids of any resolution. Where applicable, we also support situations where the input and output resolutions of the model differ, such as super-resolution.
  • Easy-to-use and beginner-friendly: Out of the box, NeuralOperator provides all necessary functionality to apply neural operator models to scientific machine learning problems in the real world. It contains simple interfaces to prebuilt neural operator architectures from the literature, subcomponent layers, a Trainer module that automates the common procedure of training a neural operator, and additional submodules, including data modules and common loss functions for training neural operators. NeuralOperator is built on top of PyTorch (Paszke et al., 2019), so we follow the torch.nn.Module convention for implementing models and their forward passes based on submodule layers.
@@
The library offers a variety of prebuilt operator layers that can be combined to build any existing or new neural operator model. These layers can be broken down by category into integral transforms (Li et al., 2020, 2021; Kovachki et al., 2023; Bonev et al., 2023), pointwise operators, positional embeddings, multi-layer blocks, and extra functionality (padding, normalization, interpolation, etc.).
@@
NeuralOperator provides state-of-the-art neural operator architectures and training recipes. End-to-end training examples to apply these architectures to various problems make it easy for newcomers to get started by adapting them to their own applications. The core building block of neural operators is the integral transform, a learnable map between two functions supplied at any two meshes. We provide several implementations for different neural operator architectures. Using a general learnable kernel integral, Graph Neural Operators (GNOs) (Li et al., 2020) allow learning on arbitrary geometries. When inputs are defined over a regular grid, the kernel integral can be efficiently realized through a Fast Fourier Transform (FFT), resulting in the spectral convolution layer used in Fourier Neural Operator (FNOs) (Li et al., 2021). We also implement Tensorized Fourier Neural Operators (TFNOs) (Kossaifi et al., 2023), which use tensor decompositions for the parameters of the spectral convolution and architectural improvements of FNOs. When working on a sphere, we can leverage the Spherical Fourier transform, which is implemented via the spherical harmonic transform as proposed in Spherical Fourier Neural Operators (SFNOs) (Bonev et al., 2023). The advantages of GNOs and FNOs can be combined using Geometry-informed Neural Operators (GINOs) (Li et al., 2023). This library also includes Risk-Controlling Neural Operators (UQNOs) (Ma et al., 2024), providing uncertainty estimates simultaneously for all points on the domain, and neural operators supplemented by locally supported kernels to learn differential and integral operators (LocalNOs) (Liu-Schiaffini et al., 2024). PDE residuals and physics losses are also supported, allowing to train and fine-tune Physics-Informed Neural Operators (PINOs) (Li et al., 2024).
@@
NeuralOperator also provides a suite of functionalities to simplify the training and deployment of neural operators. We provide a flexible DataProcessor module to pipeline all normalization and additional transformations needed to convert inputs into the form expected by the model’s forward pass and raw outputs into the form expected for computing losses. These also ensure that discretization convergence is respected in operations such as domain padding. We also provide a minimal Trainer module that automates the standard logic of a machine-learning training loop, applying the data processor, optimizing model parameters, and tracking validation metrics throughout the course of training. The Trainer is modular and highly configurable; users can provide their own models, data, and objectives, and the interface provides users with the opportunity to add custom training logic for domain-specific applications, including meta-algorithms such as incremental learning (George et al., 2024). Newcomers can directly use it on their own applications, while advanced users can directly import the neural architectures and core components and modules they need into their own workflow.

### v5

@@ Abstract
We present NeuralOperator, an open-source Python library for operator learning. Neural operators generalize neural networks to maps between function spaces instead of finite-dimensional Euclidean spaces. They can be trained and inferenced on input and output functions given at various discretizations, satisfying a discretization convergence properties. Part of the official PyTorch Ecosystem, NeuralOperator provides all the tools for training and deploying neural operator models, as well as developing new ones, in a high-quality, tested, open-source package. It combines cutting-edge models and customizability with a gentle learning curve and simple user interface for newcomers.
@@ 1 Introduction
Deep neural networks have been considered to accelerate the solution of PDEs by mapping from parameters directly to the solution on a given discretization. However, they can only learn mappings between finite-dimensional spaces, not function spaces. In other words, the solutions learned are tied to a fixed discretization. In particular, there is no guarantee that neural networks generalize to other discretizations, and they often perform poorly when interpolated to higher resolutions (Azizzadenesheli et al., 2024).
To address these limitations, a new class of machine learning models, known as neural operators, was proposed (Li et al., 2021; Kovachki et al., 2023; Berner et al., 2025). While standard neural networks learn mappings between fixed-size discretizations, i.e., finite-dimensional spaces, neural operators can directly learn mappings between functions, i.e., infinite-dimensional spaces (Kovachki et al., 2024; Berner et al., 2025). They are built from first principles to ensure a discretization convergence property: a neural operator, with a fixed set of parameters, can be applied to input functions given at any discretization. Specifically, for an input function given at various discretizations, the outputs differ only by a discretization error that converges to zero as the discretization is refined.
@@ 2 The NeuralOperator Library
NeuralOperator is an official PyTorch Ecosystem [^1] https://landscape.pytorch.org/ project, open-sourced under MIT license [^2] NeuralOperator is hosted at https://github.com/neuraloperator/neuraloperator.. The NeuralOperator library builds on the following guiding principles:
  • Resolution-agnostic: As the crucial difference to existing frameworks, modules in NeuralOperator, such as data loaders, architectures, and loss functions, should be applicable to functions at various discretization. As an example, our implementation of the Fourier Neural Operator (FNO) (Li et al., 2021; Duruisseaux et al., 2025), just like the theoretical model, can take inputs on regular grids of any resolution. Where applicable, we also support situations where the input and output resolutions of the model differ, such as super-resolution.
  • Easy-to-use and beginner-friendly: Out of the box, NeuralOperator provides all necessary functionality to apply neural operators to real-world scientific machine learning problems. It contains simple interfaces to prebuilt neural operator architectures from the literature, subcomponent layers, a Trainer module that automates the common training procedure, and additional submodules, including data modules and common loss functions for training neural operators. The library provides example training scripts and a detailed documentation, including an API reference, a user guide, and tutorials on layers and scientific datasets, from basics to advanced customization. A comprehensive practical guide to FNOs with NeuralOperator is also available (Duruisseaux et al., 2025).
@@
The library offers a variety of prebuilt operator layers, implemented following the torch.nn.Module convention, that can be combined to build any existing or new neural operator model. These layers can be broken down by category into integral transforms (Li et al., 2020, 2021; Kovachki et al., 2023; Bonev et al., 2023), pointwise operators, positional embeddings, multi-layer blocks, and extra functionality (padding, normalization, interpolation, etc.).
@@
NeuralOperator provides state-of-the-art neural operator architectures and training recipes. The core building block of neural operators is the integral transform, a learnable map between two functions supplied at any two meshes. We provide several implementations for different neural operator architectures. Using a general learnable kernel integral, Graph Neural Operators (GNOs) (Li et al., 2020) allow learning on arbitrary geometries. When inputs are defined over a regular grid, the kernel integral can be efficiently realized through a Fast Fourier Transform, resulting in the spectral convolution layer used in FNOs (Li et al., 2021). We also implement Tensor FNOs (TFNOs) (Kossaifi et al., 2023), which use low-rank tensor decompositions. When working on a sphere, we can leverage the Spherical Fourier transform, which is implemented via the spherical harmonic transform as proposed in Spherical FNOs (SFNOs) (Bonev et al., 2023). The efficiency of FNOs and SFNOs on structured grids can be combined with GNOs or optimal transport, yielding the Geometry-informed Neural Operators (GINOs) (Li et al., 2023; Lin et al., 2025) and Optimal Transport Neural Operators (OTNOs) (Li et al., 2025). Locally supported kernels to learn differential and integral operators (LocalNOs) (Liu-Schiaffini et al., 2024) can also supplement spectral convolutions. Recurrent Neural Operators (RNOs) (Liu-Schiaffini et al., 2023) extend FNOs to systems with long-term temporal dependencies by introducing recurrence and memory directly on function spaces. Physics-Informed Neural Operators (PINOs) (Li et al., 2024; Ganeshram et al., 2025) can also be trained and finetuned leveraging derivative computation routines and other functionalities available in NeuralOperator.
@@
NeuralOperator provides a suite of functionalities to simplify the training and deployment of neural operators. We provide a flexible DataProcessor to pipeline all normalization and transformations needed to convert inputs and outputs into the forms expected by the model and for computing losses. These also ensure that discretization convergence is respected in operations such as domain padding. We also provide a minimal Trainer that automates the standard logic of a machine-learning training loop, applying the data processor, optimizing model parameters, and tracking validation metrics throughout training. The Trainer is modular and highly configurable; users can provide their own models, data, and objectives, and the interface provides users with the opportunity to add custom training logic for domain-specific applications, including meta-algorithms such as incremental learning (George et al., 2024). Newcomers can directly use it on their own applications, while advanced users can directly import the neural architectures and core components and modules they need into their own workflow.